C=50+0.75Y
1. At what level of income saving become zero?
2. If autonomous investment with above consumption function is 200.at what level saving is equal to investment?
1
Expert's answer
2020-12-01T10:26:40-0500

1.

Y=G+I+C

Y=50+0.75Y

0.25Y=50


"\frac {50}{0.25}=200"

2.


"200=50+0.75Y"

"Y=\frac {150}{0.75}"


"Y=200"